OpenStudy (anonymous):

12/32 in simplest form?

OpenStudy (anonymous):

cant you divide it by 4?

OpenStudy (anonymous):

gotta get rid of the fluff... let's rewrite as products and see what we can cancel: \[\frac{ 12 }{ 32 }=\frac{ 3 \cdot \cancel4 }{ 8 \cdot \cancel4 }=\frac{ 3 }{ 8 }\]

OpenStudy (anonymous):

thank you ;D